9^x=6561 is equivalennt to?
Found 2 solutions by Cromlix, Edwin McCravy:
Answer by Cromlix(4381)

Hi there,
9^x = 6561
x log 9 = log 6561
x = log6561/log9
x = 4
Hope this helps :-)

Answer by Edwin McCravy(20054)

We shouldn't use logarithms if we can avoid it.

9x = 6561

We recognize 9 as 32

Perhaps 6561 is also a power of 3.  Let's find out:

6561/3 = 2187
2187/3 = 729
729/3 = 243
243/3 = 81
81/3 = 27
27/3 = 9
9/3 = 3
3/3 = 1

Therefore 6561 = 38 since it takes
8 divisions by 3 to get to 1.

So 

9x = 6561

(32)x = 38

32x = 38

Since the bases are the same positive number
that does not equal to 1, we may write the
eponents equal to each other:

2x = 8

 x = 4

Edwin